OpenStudy (anonymous):

Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770? I need the answer and how you got it please! :)

OpenStudy (anonymous):

x+y=150 45x+28.50y=4770 45*30=1350 (x=30) 28.5*120=3420 (y=120) 1350+3420=4770 45(30)+28.50(120)=4770

OpenStudy (anonymous):

S=120 STUDENT TICKETS WERE SOLD. 150-120=30 ADULT TICKETS WERE SOLD.